OpenStudy (anonymous):

compare each set of equations. Write a short summary of your findings for each. y=log[5]x and y=5^x

OpenStudy (saifoo.khan):

the left equation can be written as, \[y = \frac{\log x}{\log 5}\]

OpenStudy (anonymous):

So how does that compare the two though?

OpenStudy (saifoo.khan):

i dont really know.. the left one has log while the right one is more in exponential form.

OpenStudy (anonymous):

Would it be as if I could take the equation on the right, and rewrite it as a logarithm, and I could take the one on the left and rewrite it in exponential form?

OpenStudy (saifoo.khan):

kindda..

OpenStudy (anonymous):

But how could I turn the exponential form into a logarithm and how could I turn the logarithm into an exponential equation?

OpenStudy (saifoo.khan):

y=5^x this can be written as, \[x = \log_5 y\]

OpenStudy (saifoo.khan):

(if im not wrong)

OpenStudy (anonymous):

The first equation can be written as y= 5^x The second equation can be written as y= log(x)/log(5) Is that correct? or...